### v2

# No [CII] or dust detection in two Little Red Dots at

No [CII] or dust detection in two Little Red Dots at
Little Red Dots (LRDs) are compact, point-like sources characterized by their red color and broad Balmer lines, which have been debated to be either dominated by active galactic nuclei (AGN) or dusty star-forming galaxies (DSFGs). Here we report two LRDs (ID9094 and ID2756) at , recently discovered in the JWST FRESCO GOODS-North field. Both satisfy the āv-shapeā colors and compactness criteria for LRDs and are identified as Type-I AGN candidates based on their broad H emission lines (full width at half maximum: km s-1 for ID9094 and km s-1 for ID2756) and narrow [Oāiii]ā lines ( km s-1). To investigate their nature, we conduct deep NOEMA follow-up observations targeting the [Cāii] emission line and the 1.3 mm dust continuum. We do not detect [Cāii]ā or 1.3 mm continuum emission for either source. Notably, in the scenario that the two LRDs were DSFGs, we would expect significant detections: for [Cāii]Ā and for the 1.3 mm continuum of ID9094, and for [Cāii]Ā of ID2756. Using the upper limits of [Cāii]ā and 1.3 mm, we perform two analyses: (1) UV-to-FIR spectral energy distribution (SED) fitting with and without AGN components, and (2) comparison of their properties with the āSFRtot empirical relation. Both analyses are consistent with a scenario where AGN activity may contribute to the observed properties, though a dusty star-forming origin cannot be fully ruled out. Our results highlight the importance of far-infrared observations for studying LRDs, a regime that remains largely unexplored.
Key Words.:
galaxies: high-redshift ā galaxies: active ā galaxies: star-formation ā galaxies: photometry ā submillimetre: galaxies1 Introduction
Little Red Dots (LRDs; e.g., Fujimoto etĀ al., 2022; Labbe etĀ al., 2025; Matthee etĀ al., 2024; Barro etĀ al., 2024) have emerged as one of the most intriguing populations studied with the JWST (Gardner etĀ al., 2023). The term āLRDsā was first introduced by Matthee etĀ al. (2024) to describe galaxies exhibiting broad H emission lines, with a red point-source morphology. In a broader context, the name LRD has also been applied to a variety of compact, red sources ā predominantly at high redshifts () ā with āv-shapedā (blue UV and red optical slopes in rest-frame) spectral energy distributions (SEDs), based mainly on photometric measurements only. The samples of these sources overlap, contributing to the complexity of their classification. Most recent studies agree that LRDs involve a combination of stellar and AGN contributions, although there is ongoing debate over which component dominates (e.g., Wang etĀ al., 2024; Akins etĀ al., 2025a). Some works emphasize a dusty star-forming galaxy-dominated interpretation (e.g., PĆ©rez-GonzĆ”lez etĀ al., 2024; LabbĆ© etĀ al., 2023; Baggen etĀ al., 2024; Williams etĀ al., 2024), while others find stronger evidence for AGN dominance (e.g., Kocevski etĀ al., 2023, 2025; Greene etĀ al., 2024; Kokorev etĀ al., 2024; Ji etĀ al., 2025; Naidu etĀ al., 2025; de Graaff etĀ al., 2025).
Multi-wavelength studies of LRDs have primarily focused on rest-frame UV, optical, near-infrared (NIR), mid-infrared (MIR), X-ray, and radio data, whereas far-infrared (FIR) studies remain relatively scarce in comparison. FIR observations, however, are crucial for disentangling the roles of AGN activity and dusty starburst processes, as they directly probe dust properties, star formation rates, and possible AGN contributions. Some studies have started to investigate LRDs in the FIR (e.g., Labbe etĀ al., 2025; Casey etĀ al., 2024, 2025; Akins etĀ al., 2024, 2025b; Williams etĀ al., 2024; Setton etĀ al., 2025), but most rely on photometric redshifts ā either through stacked analyses or individual source studies ā which might introduce large uncertainties in interpreting their dust properties and star formation activity. Current evidence shows that many LRDs are faint in X-rays (e.g., Ananna etĀ al., 2024; Maiolino etĀ al., 2025; Yue etĀ al., 2024; Sacchi & Bogdan, 2025), exhibit weak or no radio emission (e.g., Akins etĀ al., 2024), and lack detectable hot dust in the MIR (e.g., PĆ©rez-GonzĆ”lez etĀ al., 2024; Williams etĀ al., 2024). These properties could result from heavily obscured AGNs, extreme dusty starbursts, or even hybrid scenarios. Without robust FIR constraints, however, distinguishing between these possibilities remains challenging.
Another potential issue is the partial overlap in photometric selection methods for LRDs (e.g., Labbe et al., 2025; Greene et al., 2024; Kokorev et al., 2024) and optical/NIR-selected DSFGs (optically dark/faint galaxies; e.g., Wang et al., 2019; Franco et al., 2018; Alcalde Pampliega et al., 2019; Williams et al., 2019; Gómez-Guijarro et al., 2023; Xiao et al., 2023; Gómez-Guijarro et al., 2022; Barrufet et al., 2023; Williams et al., 2024; McKinney et al., 2023; Akins et al., 2023; Pérez-GonzÔlez et al., 2023; Barro et al., 2024), due to their similar red colors. This may pose a challenge for DSFG studies, as optical/NIR-selected DSFGs typically have more relaxed selection criteria, relying solely on color and/or magnitude cuts without considering morphology. As a result, LRDs could inadvertently be included in DSFG samples. Additionally, studies also show that some DSFGs have compact morphology (e.g., Gómez-Guijarro et al., 2023). This overlap introduces ambiguity, as both populations could appear in the same sample category despite potentially distinct physical origins. For example, DSFGs are typically powered by intense star formation obscured by large amounts of dust, while AGN-driven LRDs may exhibit similar photometric properties due to their reddened SEDs but have fundamentally different drivers of energy output.
In this study, we investigate two LRDs, ID9094 and ID2756111ID9094 and ID2756 correspond to gn17537 and gn28924, respectively, in the FRESCO [Oāiii]ā catalog (Meyer etĀ al., 2024)., located in the JWST FRESCO GOODS-North field (Oesch etĀ al., 2023). Due to the overlap in photometric selection criteria between DSFGs and LRDs, these two sources were initially classified as optically faint, dust-obscured galaxy candidates with red colors (F182M F444W 1.5 magnitude; hereafter mag) and as the most massive candidates at () based on optical-to-near-IR SED fitting (Xiao etĀ al., 2024). However, as revealed in this study, deep JWST NIRCam grism spectra later reveal the presence of broad H emission lines (broad component of with full width at half maximum 1000 km s-1; e.g., Matthee etĀ al. (2024); Sect.Ā 2.4)222The broad lines were not clearly identified in the initial FRESCO spectra or at the time of the NOEMA proposal, due to the limited sensitivity of the FRESCO data., along with a red, point-source-like morphology, indicating potential AGN activity.
To further investigate their nature, we present follow-up FIR observations from NOEMA, targeting the [Cāii] emission line and the 1.3 mm dust continuum. This study, together with the recent work by Setton etĀ al. (2025), represents one of the first efforts to investigate LRDs with both dust and [Cāii]ā emission. Setton etĀ al. (2025) focused on two of the most luminous LRDs at and reported a deficit of hot and cold dust emission. In comparison, our study targets two LRDs at , a redshift regime more representative of the typical LRD population identified in JWST surveys. Together, the two studies provide a more complete view of the dust and gas properties of LRDs. Our NOEMA observations provide crucial constraints on the dust temperature, allowing us to assess whether the observed properties of these LRDs are consistent with DSFGs or if AGN-driven processes dominate. By combining FIR data with existing multi-wavelength photometry, this work addresses the critical question of whether DSFG scenarios can be ruled out for LRDs in the absence of spectroscopic confirmation.
This paper is organized as follows. In SectionĀ 2, we describe the data and observational details. SectionĀ 3 presents our results, including the non-detection of the dust continuum and [Cāii]āline in the FIR and its implications for the nature of these LRDs. In SectionĀ 4, we discuss the impact of LRDs on our understanding of DSFGs. We summarize our conclusions in SectionĀ 5.
Throughout this paper, we adopt a Chabrier initial mass function (IMF; Chabrier, 2003) to estimate SFR and stellar mass. We assume a Planck cosmology (Planck Collaboration et al., 2020) with . When necessary, data from the literature have been converted with a conversion factor of (Salpeter, 1955, IMF) = 1.7 (Chabrier, 2003, IMF). All magnitudes are in the AB system (Oke & Gunn, 1983), such that log(Sν [Jy]).
2 Data and sample
The two sources ID9094 (, [J2000] = 189.019240, 62.243531) at and ID2756 (, [J2000] = 189.083488, 62.202579) at are located in the GOODS-N field, and were first discovered in the JWST FRESCO survey (Oesch etĀ al., 2023) and reported in Xiao etĀ al. (2024). They have red colors (F182M F444W 1.5 mag) and are faint in the F182M band (F182M 26.4 mag). As such, they were selected as optically faint dusty galaxy candidates, with high inferred stellar masses. In this study, based on their broad H emission line, āv-shapeā colors, and compact morphology in F444W, we now identify both sources as LRDs, following commonly used criteria (e.g., Matthee etĀ al. 2024; Labbe etĀ al. 2025; Greene etĀ al. 2024; Kokorev etĀ al. 2024; Kocevski etĀ al. 2025; see details in Sect.Ā 2.4).
2.1 JWST observations
The two sources were initially observed by the JWST FRESCO survey (Oesch etĀ al., 2023), which provides NIRCam/grism spectroscopy in the F444W filter and direct imaging in F182M, F210M, and F444W over 62 arcmin2 in each GOODS field, North and South. The grism spectra cover a wavelength range of 3.8 to 5.0 m at a resolution of R1,600, reaching a 5 line sensitivity of 10-18 ergās-1ācm-2. The images reach typical 5 depths of 28.3, 28.1, and 28.2 mag, respectively, in a 016 circular aperture. Data reduction was performed using the grizli pipeline333https://github.com/gbrammer/grizli, including continuum subtraction and optimal 1D spectral extraction.
Recently, we obtained deeper NIRCam/grism spectroscopy from a Cycle 3 JWST program (GO-4762; PIs: S. Fujimoto & G.Ā Brammer), using the F410M filter instead of F444W. This choice improves spectral sensitivity by blocking the high thermal background beyond 4.4 m, at the cost of reduced wavelength coverage. Despite a slightly shorter exposure time (by a factor of 1.3) than FRESCO, this dataset thus achieves a 5 line sensitivity of 10-19 ergās-1ācm-2. The continuum subtraction was optimized to explicitly mask emission lines and a wide kernel with a large central hole (151, 31 pixels, respectively; for the running median) to prevent over-subtraction of broad wings (see Matthee etĀ al., 2024).
The NIRCam images and grism spectra of ID9094 and ID2756 are shown in Fig.Ā 1, updated with the latest data.
| Instrument | Units | ID2756 | ID9094 | |
| Chandraa | 0.5ā7 keV | ergās-1ācm-2 | ||
| 2ā7 keV | ergās-1ācm-2 | |||
| 0.5ā2 keV | ergās-1ācm-2 | |||
| HST/ACS F435W | 0.43āµm | nJy | ||
| HST/ACS F606W | 0.59āµm | nJy | ||
| HST/ACS F775W | 0.77āµm | nJy | ||
| HST/ACS F814W | 0.80āµm | nJy | ||
| HST/ACS F850LP | 0.90āµm | nJy | ||
| HST/WFC3 F105W | 1.06āµm | nJy | 33.0 8.4 | ⦠|
| HST/WFC3 F125W | 1.25āµm | nJy | 87.5 10.8 | 75.7 9.8 |
| HST/WFC3 F140W | 1.40āµm | nJy | 84.7 6.6 | ⦠|
| HST/WFC3 F160W | 1.54āµm | nJy | 91.3 7.8 | 109.7 10.4 |
| JWST/NIRCam F090W | 0.90āµm | nJy | ||
| JWST/NIRCam F115W | 1.15āµm | nJy | 79.2 4.4 | 64.8 6.2 |
| JWST/NIRCam F150W | 1.50āµm | nJy | ⦠| 93.7 12.9 |
| JWST/NIRCam F182M | 1.85āµm | nJy | 80.0 4.0 | 99.3 6.7 |
| JWST/NIRCam F200W | 1.99āµm | nJy | ⦠| 109.6 11.5 |
| JWST/NIRCam F210M | 2.10āµm | nJy | 117.8 5.9 | 130.2 8.8 |
| JWST/NIRCam F356W | 3.57āµm | nJy | 225.8 11.3 | 443.8 22.2 |
| JWST/NIRCam F444W | 4.40āµm | nJy | 372.4 18.6 | 887.7 44.4 |
| Spitzer/IRAC | 3.6āµm | nJy | 191.0 15.4 | 400.4 13.0 |
| 4.5āµm | nJy | 340.2 17.6 | 592.7 16.5 | |
| 5.8āµm | nJy | 933.7 243.5 | ||
| 8.0āµm | nJy | 351.2 284.8 | 2683.0 289.4 | |
| Spitzer/MIPSb | 24āµm | µJy | ||
| 70āµm | mJy | |||
| Herschel/PACSb | 100āµm | mJy | ||
| 160āµm | mJy | |||
| Herschel/SPIREb | 250āµm | mJy | ||
| 350āµm | mJy | |||
| 500āµm | mJy | |||
| SCUBA-2c | 850āµm | mJy | ||
| NOEMA | 1.3āmm | µJy | ||
| [Cāii]158μm | Jy km s-1 | |||
| NIKA-2d | 1.2āmm | µJy | ||
| 2.0āmm | µJy | |||
| VLA 1.5āGHze | 20ācm | uJy |
ā¢
Note: We use 1 for uncertainties and 3 upper limits for non-detections. (a) We calculate 3 upper limits based on the mean sensitivity limits achieved in the central 1 arcmin2 area at the average aim point from 2 Ms Chandra Deep Field-North survey (Xue etĀ al., 2016). (b) Sources are not detected, we adopt 3 upper limits from the GOODS-Herschel survey (Elbaz etĀ al., 2011). (c) We adopt the confusion limit for a 3 detection in the central region of GOODS-N from Cowie etĀ al. (2017). (d) We adopt 3 upper limits according to the N2CLS depth from Bing etĀ al. (2023). (e) Calculated based on the rms noise for a 16 resolution in the field center from Owen (2018).
| ID | FluxHβ,broad | FluxHβ,narrow | Flux | |||||
| erg s-1cm-2 | erg s-1cm-2 | erg s-1cm-2 | km s-1 | km s-1 | ||||
| ID9094 | ||||||||
| ID2756 |
ā¢
Note. and are continuum slopes derived from multi-band JWST photometry blueward and redward of the Balmer break, respectively, and trace the two sides of the āv-shapeā SEDs (see Sect.Ā 2.4).
2.2 NOEMA observations
The NOEMA observations of ID9094 and ID2756 were carried out in the summer of 2023 using the PolyFiX correlator (Project ID: S23CY, PIs: D. Elbaz & M. Xiao). Conducted with array configuration D in band 3, the observations targeted the [Cāii] emission line, and the 1.3 mm dust continuum, with representative frequencies of 232.052 GHz for ID9094 and 236.576 GHz for ID2756. The total on-source time was 6 hours and 4.5 hours for ID2756 and ID9094, respectively.
All the calibrations and the creation of table were performed using CLIC package from the IRAM GILDAS software444https://www.iram.fr/IRAMFR/GILDAS/ with the support of IRAM astronomers. We then performed further analysis with version 6.5.5 of the Common Astronomy Software Application package (CASA; McMullin etĀ al., 2007).
For the [Cāii] emission line, imaging was carried out using the tclean task with 0.4ā²ā² pixels and a channel width of 50 km s-1 with a natural weighting. The natural weighting provides the best point-source sensitivity, which is optimal for source detection. The resulting data cube has a synthesized beam size of full width at half maximum (FWHM) 1838 1472 (9.8 kpc 7.8 kpc in physical scale) and 2859 1854 ( 15.0 kpc 9.7 kpc) with a root mean square (rms) sensitivity of 0.45 and 0.84 mJy beam-1 per channel at the phase center, for ID9094 and ID2756, respectively.
We also created the observed 1.3 mm continuum maps using natural weighting and the same method as for the [Cāii] emission line. For ID9094, the rms level is 30.4 Jy beam-1 in the map of 1725 1383 ( 9.2 kpc 7.4 kpc) angular resolution. For ID2756, the rms level is 50.3 Jy beam-1 in the map of 2712 1743 ( 14.3 kpc 9.2 kpc) angular resolution.
2.3 Multi-wavelength observations
Besides the JWST FRESCO and NOEMA observations, the two sources are also covered by a wide array of multi-wavelength observations. In Table 1, we summarize the multi-wavelength dataset for our two sources from X-ray to radio: () X-ray: Chandra 2 Ms (0.5ā7.0keV, 0.5ā2.0keV, and 2ā7keV bands) images in the Chandra Deep Field-North field (CDF-N; Xue etĀ al., 2016); () HST/ACS (F435W, F606W, F775W, F814W, F850LP) and HST/WFC3 (F105W, F125W, F140W, F160W) images from the Hubble Legacy Fields Program (HLF555https://archive.stsci.edu/prepds/hlf/; Whitaker etĀ al., 2019); () JWST/NIRCam (F090W, F115W, F150W, F182M, F200W, F210M, F356W, F444W) from FRESCO (Oesch etĀ al., 2023), CONGRESS (Egami etĀ al., 2023), and JADES (Eisenstein etĀ al., 2023; DāEugenio etĀ al., 2025). The two sources are not yet covered by any publicly available JWST/MIRI observations; () Spitzer/IRAC from the GREATS program (3.6 āµm, 4.5 āµm, 5.8 āµm, 8 āµm; Stefanon etĀ al., 2021, new catalogs based on FRESCO sources are taken from Stefanon et al. in prep; priv. comm.); () Spitzer/MIPS, Herschel/PACS and SPIRE catalog is taken from the GOODS-Herschel survey (Elbaz etĀ al., 2011); () JCMT/SCUBA-2 at 850 āµm (Cowie etĀ al., 2017); () NOEMA data from this study (see Sect. 2.2); () NIKA2 Cosmological Legacy Survey at 1.2 mm and 2.0 mm (Bing etĀ al., 2023); and () VLA 1.5 GHz (Owen, 2018).
We note that the photometric measurements from HST and JWST images are derived following the same procedure as outlined in Weibel etĀ al. (2024). Briefly, we use SExtractor (Bertin & Arnouts, 1996) in dual image mode with an inverse-variance weighted stack of the F210M+F444W with FRESCO-only imaging, as the detection image. In this study, fluxes are measured in 016 radius circular apertures in images that are point spread function (PSF)-matched to the F444W band. Total fluxes are derived from the Kron AUTO aperture provided by SExtractor in the F444W band, in addition to a correction based on the encircled energy of the Kron aperture on the F444W PSF. Detailed descriptions of data reduction and photometric measurements are provided in Weibel etĀ al. (2024) and Xiao etĀ al. (2024).
2.4 Our sample: two āLittle Red Dotsā at
Although there is no uniform definition of LRDs, we identify our two sources as LRDs because they satisfy both of two commonly used selection criteria: 1) an initial criterion ā sources with broad Balmer emission lines ( km s-1; Matthee etĀ al., 2024); and 2) a widely used criterion based only on photometric measurements ā a combination of āv-shapeā SED and compactness selection (e.g., Labbe etĀ al., 2025; Kokorev etĀ al., 2024; Greene etĀ al., 2024; Kocevski etĀ al., 2025).
Fig.Ā 1 presents the JWST/NIRCam images and grism spectra of ID9094 and ID2756. The RGB images show that both sources appear red with point-like morphology. The grism spectra show H and [Oāiii] emission lines. Both sources exhibit a broad H line. We fit the H line with both narrow and broad Gaussian components, assuming the narrow component has the same width as the [Oāiii] lines. The best-fit line properties are shown in TableĀ 2. The broad component of H has a FWHM of km s-1 for ID9094 and km s-1 for ID2756, suggesting a possible Type I (broad emission line) AGN. Notably, ID2756 has also been identified as a strong Lyman- emitter, with a considerable equivalent width of 221 Ć Ā (Leonova et al. in prep.).
We further compare our two sources to an alternative color and compactness selection criterion for LRDs, designed to identify compact, red objects with āv-shapeā SEDs (e.g., Kokorev etĀ al., 2024; Greene etĀ al., 2024):
Both sources are bright in F444W, with magnitudes of mag (ID9094) and mag (ID2756), consistent with the LRD selection. Since ID2756 lacks observations in F277W and F200W, while ID9094 is missing F277W, the corresponding fluxes used here are derived from the best-fit SED and the 16thā84th percentile range (as shown later in Fig.Ā 4). The detailed comparison is shown in Fig.Ā 2. ID9094 lies within the red region, perfectly matching the LRD selection. ID2756 exhibits a āv-shapedā color and compact morphology, aligning with the LRD classification. Its F277W ā F356W color is slightly bluer than the selection threshold but still within the 1 uncertainty. We therefore classify both sources as LRDs. Compared to the majority of sources from various JWST blank legacy fields (blue-shaded region in Fig.Ā 2), obtained from the DAWN JWST Archive (DJA666https://dawn-cph.github.io/dja/imaging/v7/; Valentino etĀ al., 2023), these two LRDs stand out as particularly exceptional, exhibiting redder colors and more compact morphologies.
To further quantify the āv-shapeā SEDs, we also measured the UV and optical continuum slopes ( and ) using multi-band photometry, following the method of Kocevski etĀ al. (2025). Both sources satisfy their LRD selection criteria ( and ). The measured slope values are listed in TableĀ 2. In addition, both sources are unresolved in F444W, further supporting that they are typical LRDs.
Altogether, we find that both sources are consistent with standard criteria for LRDs based on their morphology and āv-shapeā SED. In addition, both sources show the broad H line, making them strong Type I AGN candidates.
3 Results
3.1 Non-detections in both [Cāii] and 1.3 mm of two LRDs
We analyze the NOEMA data cubes and extract spectra at the positions of our sources using an aperture consistent with the beam size. The resulting spectra, shown in Fig.Ā 3, reveal no significant signal at the expected [Cāii]Ā frequency (). Assuming a [Cāii] line width of 250 km s-1 ā a typical value for high-redshift galaxies (BĆ©thermin etĀ al., 2020; Endsley etĀ al., 2022) ā we construct the velocity-integrated intensity map (moment-0). Fig.Ā 3 overlays the contours from the [Cāii]Ā moment-0 and 1.3 mm maps on the JWST RGB image, showing no detections () at the source positions and in the surrounding regions. We therefore conclude that neither the [Cāii] emission line nor the 1.3 mm dust continuum emission are detected for the two LRDs in the current NOEMA data.
We then determine 3 upper limits for both [Cāii]Ā and dust continuum. Using the rms of the 1.3 mm map (see Sect.Ā 2.2) and assuming our sources are not resolved, we derive 3 upper limits of 91 Jy and 151 Jy for ID9094 and ID2756, respectively. In addition, the 3 upper limits for [Cāii]Ā are calculated as three times the pixel-by-pixel rms of the moment-0 map, which gives upper limits of 0.157 Jy km s-1 and 0.310 Jy km s-1 for ID9094 and ID2756, respectively. The corresponding 3 upper limits on the [Cāii]ā emission-line luminosity are log( for ID9094 and log( for ID2756. Using , we further calculate the 3 upper limits on the molecular gas mass (Zanella etĀ al., 2018) to be log( for ID9094 and log( for ID2756.
Intriguingly, when requesting the NOEMA observations, we assumed these two galaxies to be DSFGs. Given their stellar masses from Xiao etĀ al. (2024), we estimated their total star formation rates (SFRtot) assuming that they lie on the star-forming main sequence (SFMS; Schreiber etĀ al., 2015) and predicted their expected [Cāii]Ā fluxes using the āSFRtot relation for SFGs (Schaerer etĀ al., 2020). Similarly, we predicted their 1.3 mm continuum fluxes by assuming an infrared SED shape consistent with typical dust-obscured galaxies (optically dark/faint galaxies; Xiao etĀ al., 2023). Given the depth of our NOEMA observations, we expected ID9094 to be detected at in [Cāii]Ā and in the 1.3 mm continuum, while ID2756 should be detected in [Cāii]Ā at . However, the complete absence of [Cāii]Ā and 1.3 mm emission suggests that these two LRDs are not the typical dusty SFGs initially assumed. This result aligns with our findings in Sect.Ā 2.4, where both sources exhibit a broad H emission line, being Type-I AGN candidates.
3.2 The hints of AGNs in two LRDs from FIR non-detections
In this section, we utilize the 3 upper limits on the [Cāii] emission line and the 1.3 mm continuum to investigate the possible presence of AGN activity in the LRDs. While both sources have already been identified as strong Type-I AGN candidates based on their broad H emission lines, our primary goal here is to investigate whether the FIR non-detections can effectively rule out the DSFG scenario for LRDs, independently of spectral information. Here, we focus primarily on ID9094, which benefits from deeper NOEMA observations compared to ID2756, allowing for more robust constraints. We assess the AGN contribution through two independent approaches: 1) UV-to-FIR spectral energy distribution (SED) fitting analysis and 2) the āSFRtot empirical relation.
3.2.1 SED fitting analysis
We perform a UV-to-FIR SED fitting analysis using CIGALE (Boquien etĀ al., 2019), which is based on an energy balance principle, fixing the redshift to . We assume a constant star formation history (SFH) and adopt Bruzual & Charlot (2003) stellar population models with the Calzetti etĀ al. (2000) dust attenuation law. We adopt a broad metallicity range of 0.004 to 0.02, dust attenuation to the rest-frame V band () values from 0 to 6 magnitudes, and ionization parameters log of -2.0. The AGN component is modeled using the Fritz etĀ al. (2006) template, with the AGN fraction (), i.e., the contribution of the AGN to the total IR luminosity, ranging from 0.01 to 0.8. The angle between the equatorial axis and the viewing angle at the line of sight () ranges from 40 to 90. With a fixed open angle of the AGN torus at 100 degrees, this range of refers to Type I AGN. We note that recent MIR observations (Wang etĀ al., 2025; Williams etĀ al., 2024; Barro etĀ al., 2024) suggest a lack of hot torus dust emission in LRDs, which may deviate from the Fritz etĀ al. (2006) template adopted. However, with the absence of deep MIR observations, we anticipate little change in the quality of SED fitting on our sample by only changing the AGN torus emission in the templates. On the contrary, the inclusion of continuum from accretion disk by limiting to Type I AGN is critical, as we robustly detected broad lines in H.
Given that we only have upper limits in the FIR, we adopt a simple FIR dust model from Casey (2012), using the default dust emissivity index () and MIR power-law slope (). Compared to , which is adopted in some studies of dusty star-forming galaxies (e.g., Hildebrand, 1983; KovÔcs et al., 2006; Gordon et al., 2010; Gómez-Guijarro et al., 2022; Xiao et al., 2023), our choice of is more conservative and leads to slightly lower inferred dust temperatures for a given FIR slope. During the fitting process, we find that when the dust temperature () is allowed to vary across a defined range, the best-fit is strongly influenced by the upper boundary of the input range. This effect arises due to the absence of deep MIR data, which prevents us from constraining the maximum . Consequently, we cannot determine a precise best-fit value for , but the deep NOEMA data allow us to place a lower limit.
To refine our constraints, we adopt an optimized fitting approach that allows us to constrain the minimum allowed using the NOEMA 1.3 mm 3 upper limits. In this approach, we fix during the fitting and perform multiple fits with different values, ranging from 30 K to 150 K. In addition, we also correct the effect of the cosmic microwave background (CMB) as an observing background on photometry, following da Cunha etĀ al. (2013). The CMB has the effect of raising the apparent dust temperature with increasing redshift, as the CMB temperature scales with redshift and provides an additional background source of radiation for dust heating. We assume , consistent with previous studies for galaxies (Sommovigo etĀ al., 2022). In this way, we derive the intrinsic FIR photometry starting from MIPS 70m, which is used for SED fitting. The best-fit SEDs, with and without AGN components considered, are presented in Fig.Ā 4.
For ID9094, without including an AGN component (Fig.4-), we present three representative examples of best-fit SEDs with () K, () K, and () K. These correspond to results that the predicted flux density at 1.3 mm exceeds, matches, and falls below the observed 3 upper limits, respectively, under the energy balance assumption of CIGALE. At K ā a typical value for star-forming UV-bright galaxies in the REBELS survey at (Sommovigo etĀ al., 2022) ā the predicted 1.3 mm flux exceeds the observed 3 upper limit by nearly a factor of three. This discrepancy suggests that, with our current NOEMA observations, we should have detected the source at a significance level of , which is clearly not the case. As increases, the peak of the FIR SED moves towards the shorter wavelength and the fit improves. At K (Fig.4-), the predicted 1.3 mm flux aligns with the observed 3 upper limit. However, given that the actual signal-to-noise ratio at 1.3 mm is well below 3, this implies that the real would be greater than 110 K.
Therefore, without including an AGN component, we obtain a lower limit of K. At fixed K, as in TableĀ 3, the best-fit IR luminosity is , yr-1 (we use this for the analysis in Sect.Ā 3.2.2), and stellar mass of . Remarkably, even this lower limit is significantly higher than the typical K observed in REBELS galaxies at similar redshifts (Sommovigo etĀ al., 2022). Similar findings have been reported in studies of 675 LRDs at , regardless of heating mechanism, whether AGN or star formation, where an average luminosity-weighted dust temperature of K is reported (Casey etĀ al., 2024). Such an extreme dust temperature is challenging to explain solely through heating by star formation. In typical star-forming galaxies, the dust temperature is primarily driven by the ultraviolet radiation field of young stars, which is usually insufficient to reach the high values observed here (see e.g., Harshan etĀ al., 2024; Laporte etĀ al., 2017).
Based on the best-fit without AGN at fixed K, we derived the dust emitting region size pc, based on a black-body assumption777 and is the StefanāBoltzmann constant.. A crucial parameter to consider is the stellar surface density, given by pc-2 (similar findings are reported in, e.g., Baggen etĀ al., 2024). This is about one order of magnitude higher than the densest systems in the local universe (Hopkins etĀ al., 2010) and SFGs at observed by JWST (Morishita etĀ al., 2024; Schaerer etĀ al., 2024), highlighting a general issue for LRD models that rely solely on stellar components. Expanding the star formation region would reduce , but it would also decrease below the observed values (e.g., De Rossi etĀ al., 2018), leading to an inconsistency with our findings. These values are difficult to explain through star formation alone under standard assumptions and may suggest the contribution of additional heating sources, such as AGN activity. However, we emphasize that this interpretation remains model-dependent, and we further discuss such caveats in Sect.Ā 3.2.3.
| ID9094 | fixed [K] | log ] | SFR [yr-1] | [] |
| Without AGN | 110 | |||
| With AGN | 60 | |||
| With AGN | 40 |
When the AGN component is included in the SED fitting (Fig.Ā 4-), the results become more physically plausible. Similar to Fig.4-, we also show three representative examples of best-fit SEDs with () K, () K, and () K. However, due to the degeneracy between the dust temperature and the AGN fraction, the current upper limit at 1.3 mm does not allow us to place a meaningful lower limit on . The observed photometry is consistent with SED fits with K and even marginally consistent with down to 40 K, a value comparable to those observed in AGN and/or quasar host galaxies at ( K; e.g., Walter etĀ al., 2022; Tsukui etĀ al., 2023; Tripodi etĀ al., 2023; Decarli etĀ al., 2023; Meyer etĀ al., 2025). At fixed K, the best-fit AGN fraction is , indicating a scenario where both AGN and star formation contribute significantly to the FIR emission. The corresponding best-fit , , and are listed in TableĀ 3. We also note that both sources are covered by the MEOW survey (JWST Cycle 3 GO-5407; PI: G. Leung; priv. comm.), which detects both in F1000W and ID9094 also in F2100W. Although the fluxes are not yet publicly released, preliminary analysis suggests that ID9094 exhibits significant hot dust emission in the MIR (G. Leung, priv. comm.).
For ID2756, we perform SED fitting using the same strategy as for ID9094 (see AppendixĀ A). However, the current 1.3 mm observations for ID2756 (rms level 50.3 Jy beam-1) are not sufficiently deep to constrain the SED effectively. The lack of sensitivity prevents us from distinguishing between the DSFG and AGN scenarios based solely on the SED fitting results. To achieve meaningful constraints, deeper FIR/submillimeter observations are required. Specifically, observations at least as deep as for ID9094 (rms level 30.4 Jy beam-1) are necessary to provide tighter limits on the 1.3 mm flux density and help refine the SED fitting.
3.2.2 āSFRtot relation
In this section, we compare the locations of our LRDs with the well-established āSFRtot scaling relation to investigate their alignment with expectations for star-forming galaxies or the potential influence of AGN activity. The āSFRtot relation has been shown to hold across a wide range of redshifts, with no significant evolution observed up to (e.g., Lagache etĀ al., 2018; Schaerer etĀ al., 2020; Carniani etĀ al., 2020; Pallottini etĀ al., 2022). Additionally, studies indicate that the relation is consistent for both AGN-hosting galaxies and typical star-forming galaxies, suggesting that the presence of AGN does not significantly alter the global āSFRtot relation (e.g., Herrera-Camus etĀ al., 2018).
For ID9094, when no AGN component is included in the SED fitting, the SFR derived corresponds to K. Combining this SFR with the 3 upper limits of the (log(; Sect.Ā 3.1) places ID9094 significantly below the āSFRtot relation, with a deviation of dex (see Fig.Ā 5). This large offset suggests that the observed properties of ID9094 may not be fully explained by typical star formation processes in dusty star-forming galaxies. In contrast, when including an AGN component in the SED fitting, the resulting SFR with the K template is about four times lower than that returned by SED fitting without an AGN component. In this case, ID9094 lies within the scatter of the āSFRtot relation, indicating consistency with the scaling relation. We note that including an AGN component with K results in a lower SFR and brings ID9094 into even better agreement with the āSFRtot relation.
For ID2756, its 3 upper limits of the [Cāii]ā emission line luminosity is log( (Sect.Ā 3.1). Due to the shallower NOEMA observations, we cannot constrain SFR well with and without AGN components in the SED (as described in Sect.Ā 3.2.1). Therefore, we are unable to place meaningful constraints on its position in the versus SFRtot plane without deeper FIR observations.
3.2.3 Caveats
Our analysis provides constraints on the dust and gas properties of the two LRDs. For ID9094, the two independent analyses of and the āSFRtot relation show notable deviations from typical dusty star-forming galaxies. These distinct properties suggest that additional radiation mechanisms, such as AGN activity, may contribute to the total emission of this source.
However, we emphasize that the interpretations above rely on several assumptions and modeling choices. First, the lower limit on the dust temperature (āK) inferred in the absence of an AGN component is derived under fixed SED parameters and based solely on upper limits in the FIR. The resulting value is therefore model-dependent and sensitive to assumptions such as the dust emissivity index and IR SED shape. Moreover, high dust temperatures alone are not sufficient to definitively rule out a purely star-forming scenario. Although rare, some high-redshift galaxies without clear AGN signatures have exhibited elevated values. For instance, MACS0416-Y1 at has a K, potentially due to intense starburst activity in a merger-driven system without requiring AGN heating (e.g., Bakx etĀ al., 2020; Fudamoto etĀ al., 2023; Harshan etĀ al., 2024; Sommovigo etĀ al., 2022). Similarly, A2744-YD4 at shows K, which may be attributed to its low metallicity and high star formation rate (e.g., Laporte etĀ al., 2017; Behrens etĀ al., 2018; Sommovigo etĀ al., 2022). These cases are considered extreme and likely involve compact starbursts and low dust content.
Second, the observed offset from the empirical āSFR relation is not conclusive evidence against the DSFG scenario. The total SFR estimates are subject to uncertainties, especially when derived from upper limits, and the comparison sample includes a mix of UV-bright and IR-selected galaxies with diverse properties. Moreover, the āSFR relation exhibits large scatter at high redshift, and many star-forming galaxies also lie below the relation.
In summary, while our results are more consistent with an AGN-driven scenario, the possibility of dusty star formation remains viable given the current data limitations. Future deeper FIR observations will be essential to distinguish between these scenarios.
4 Impact of LRDs on our understanding of DSFGs
As introduced earlier, the two LRDs analyzed in this study were initially selected as the most massive DSFG candidates at in the JWST FRESCO GOODS-North field (Xiao etĀ al., 2024). This misclassification arose due to the overlap in selection criteria between LRDs and DSFGs, particularly for optical/NIR-selected DSFGs ā often referred to as optically dark/faint galaxies (also known as HST-dark galaxies or H-dropouts; e.g., Wang etĀ al., 2019; Alcalde Pampliega etĀ al., 2019; Williams etĀ al., 2019; Xiao etĀ al., 2023; Gómez-Guijarro etĀ al., 2022; Barrufet etĀ al., 2023; Akins etĀ al., 2023; PĆ©rez-GonzĆ”lez etĀ al., 2023; Barro etĀ al., 2024). Typically, DSFGs identified in optical/NIR observations are selected based on their red colors and rest-frame optical faintness, as these characteristics correlate strongly with dust attenuation at similar redshifts (e.g., Wang etĀ al., 2016; Xiao etĀ al., 2023). Additionally, the optical faintness criterion helps exclude quiescent galaxies, which are generally optically bright. For instance, DSFGs are often selected using a color cut such as F150WāF444W mag and F150W mag, though specific band choices and thresholds slightly vary across studies.
With the increasing availability of JWST spectroscopic data, emission-line detections have become an additional criterion for DSFG selection (e.g., Xiao etĀ al., 2024; Barrufet etĀ al., 2025). In Xiao etĀ al. (2024), DSFG candidates were selected based on their red color (F182M F444W Āæ 1.5 mag) and strong emission lines (e.g., H+[NII]+[SII] lines or [Oāiii]+H lines). This approach also allows for the identification of LRDs among DSFG candidates. Among 26 DSFG candidates with detected H lines (), seven () were found to be LRDs, exhibiting broad H emission lines (see Matthee etĀ al., 2024).
Over the past decade, numerous studies have shown that optically-dark/faint galaxies tend to have high stellar masses and may dominate the massive end of the stellar mass function (e.g., Wang etĀ al., 2019; Gottumukkala etĀ al., 2024). Additionally, these galaxies could contribute significantly to the cosmic star formation rate density (SFRD) at , with contributions of of the SFRD, depending on the methodology (see a direct comparison in Xiao etĀ al. 2023; e.g., Wang etĀ al. 2019; Williams etĀ al. 2019; Gruppioni etĀ al. 2020; Fudamoto etĀ al. 2021; Talia etĀ al. 2021; Enia etĀ al. 2022; Shu etĀ al. 2022; Barrufet etĀ al. 2023; van der Vlugt etĀ al. 2023). Notably, while different studies employ varying selection criteria for optically dark/faint galaxies, some adopt stricter definitions, requiring FIR/submm (e.g., ALMA/NOEMA) detection and/or non-detection in HST. Given that LRDs have been found to 1) lack FIR/submm detections (see also e.g., Labbe etĀ al., 2025; Williams etĀ al., 2024), 2) exhibit a characteristic āv-shapedā SED (indicating detection in HST), they are unlikely to contaminate the DSFG sample in these studies.
On a global scale, these findings suggest that LRD contamination should not significantly bias SFRD estimates derived from DSFG samples. However, their presence may still impact number density estimates of optical/NIR-selected DSFG populations, particularly when broader color-based selection criteria are used. A study with a larger sample is needed to determine how much the number density estimates of DSFG populations are affected by LRDs. These results highlight the importance of refining DSFG selection methods to minimize potential contamination from AGN-dominated LRDs, ensuring a more accurate census of dust-obscured star formation in the early Universe.
5 Conclusions
In this paper, we present a study of two LRDs, ID9094 and ID2756, recently discovered in the JWST FRESCO GOODS-North field at . These sources, characterized by their red, point-source-like morphology and āv-shapedā SEDs, exhibit broad H emission lines, suggesting them as AGN candidates. Using NOEMA observations targeting the [Cāii] emission line, and the 1.3 mm dust continuum, we obtain the following key results:
ā¢
Both sources are undetected in the [Cāii]Ā and 1.3 mm continuum. For ID9094, the upper limits on [Cāii]Ā and 1.3 mm continuum are far below expectations for a typical DSFG, where it would have been detected at in [Cāii]Ā and at 1.3 mm. For ID2756, the expected detection in [Cāii]Ā is also not observed.
ā¢
Benefiting from the 1.3 mm upper limits, we perform UV-to-FIR SED fitting, which provides the first hint of an AGN scenario in this work. For ID9094, without including an AGN component, the SED fit requires an abnormally high dust temperature ( K), significantly above typical values for SFGs at similar redshifts. Including an AGN component yields more physically plausible fits with dust temperatures that can be as low as 40ā60 K. However, we caution that the minimum is strongly linked to the AGN fraction contributing to the total IR luminosity ā higher AGN fractions allow for lower values. For ID2756, the SED fitting remains inconclusive due to shallower NOEMA data, highlighting the need for deeper FIR/submillimeter observations.
ā¢
Using the [Cāii]Ā upper limits, we compare the sources to the empirical āSFRtot relation, providing a second hint of an AGN scenario in this work. For ID9094, its [Cāii]Ā upper limit places it significantly below the scaling relation when no AGN component is included, but consistent with the relation when an AGN contribution is considered. The case for ID2756 remains uncertain due to limited NOEMA depth.
These results highlight the value of FIR constraints, when combined with shorter-wavelength photometry, in revealing the nature of LRDs ā particularly when spectroscopic data are unavailable. For ID9094, the results are consistent with a scenario in which AGN activity contributes to the observed emission, though a dusty star-forming origin cannot be fully excluded given the data limitations. We also note that both sources are detected in JWST/MIRI F1000W, and ID9094 is additionally detected in F2100W, supporting the presence of hot dust emission (G. Leung, priv. comm.). For ID2756, the current observations are insufficient to determine the dominant energy source. We emphasize that our interpretation remains model-dependent and subject to assumptions in the SED fitting and limited FIR constraints. More broadly, this study demonstrates the critical role of FIR data in clarifying the nature of LRDs. Future deeper FIR/submillimeter campaigns, along with high-resolution spectroscopy and multi-wavelength data, will be essential to fully understand the physical properties of these LRDs and their role in early galaxy evolution.
Acknowledgements.
We are very grateful to the anonymous referee for instructive comments, which helped improve the overall quality and strengthen the analysis of this work. We thank Andrea Weibel for assistance with the HST and JWST photometric measurements used in this paper. This work is based on observations carried out under project number S23CY with the IRAM NOEMA Interferometer. IRAM is supported by INSU/CNRS (France), MPG (Germany) and IGN (Spain). This work is based in part on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs #1895 and #4762. Support for programs #1895 and #4762 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. This work has received funding from the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00072, as well as from the Swiss National Science Foundation (SNSF) through project grant 200020_207349. The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant DNRF140. This work is partially supported from the National Natural Science Foundation of China (12073003, 11721303, 11991052), and the China Manned Space Project (CMS-CSST-2021-A04 and CMS-CSST-2021-A06). Y.F.Ā is supported by JSPS KAKENHI Grant Numbers JP22K21349 and JP23K13149. M.V.Ā gratefully acknowledges financial support from the Independent Research Fund Denmark via grant numbers DFF 8021-00130 and 3103-00146 and from the Carlsberg Foundation via grant CF23-0417. VK acknowledges support from the University of Texas at Austin Cosmic Frontier Center. S.F. acknowledges support from NASA through the NASA Hubble Fellowship grant HST-HF2-51505.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. Support for this work for RPN was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51515.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555.References
- Akins etĀ al. (2023) Akins, H.Ā B., Casey, C.Ā M., Allen, N., etĀ al. 2023, ApJ, 956, 61
- Akins etĀ al. (2025a) Akins, H.Ā B., Casey, C.Ā M., Berg, D.Ā A., etĀ al. 2025a, ApJ, 980, L29
- Akins etĀ al. (2025b) Akins, H.Ā B., Casey, C.Ā M., Chisholm, J., etĀ al. 2025b, submitted to ApJL, arXiv:2503.00998
- Akins etĀ al. (2024) Akins, H.Ā B., Casey, C.Ā M., Lambrides, E., etĀ al. 2024, submitted to ApJ, arXiv:2406.10341
- Alcalde Pampliega et al. (2019) Alcalde Pampliega, B., Pérez-GonzÔlez, P. G., Barro, G., et al. 2019, ApJ, 876, 135
- Ananna etĀ al. (2024) Ananna, T.Ā T., BogdĆ”n, Ć., KovĆ”cs, O.Ā E., Natarajan, P., & Hickox, R.Ā C. 2024, ApJ, 969, L18
- Baggen etĀ al. (2024) Baggen, J. F.Ā W., van Dokkum, P., Brammer, G., etĀ al. 2024, ApJ, 977, L13
- Bakx etĀ al. (2020) Bakx, T. J.Ā L.Ā C., Tamura, Y., Hashimoto, T., etĀ al. 2020, MNRAS, 493, 4294
- Barro et al. (2024) Barro, G., Pérez-GonzÔlez, P. G., Kocevski, D. D., et al. 2024, ApJ, 963, 128
- Barrufet etĀ al. (2025) Barrufet, L., Oesch, P.Ā A., Marques-Chaves, R., etĀ al. 2025, MNRAS, 537, 3453
- Barrufet etĀ al. (2023) Barrufet, L., Oesch, P.Ā A., Weibel, A., etĀ al. 2023, MNRAS, 522, 449
- Behrens etĀ al. (2018) Behrens, C., Pallottini, A., Ferrara, A., Gallerani, S., & Vallini, L. 2018, MNRAS, 477, 552
- Bertin & Arnouts (1996) Bertin, E. & Arnouts, S. 1996, A&AS, 117, 393
- Béthermin et al. (2020) Béthermin, M., Fudamoto, Y., Ginolfi, M., et al. 2020, A&A, 643, A2
- Bing et al. (2023) Bing, L., Béthermin, M., Lagache, G., et al. 2023, A&A, 677, A66
- Boquien etĀ al. (2019) Boquien, M., Burgarella, D., Roehlly, Y., etĀ al. 2019, A&A, 622, A103
- Bruzual & Charlot (2003) Bruzual, G. & Charlot, S. 2003, MNRAS, 344, 1000
- Calzetti etĀ al. (2000) Calzetti, D., Armus, L., Bohlin, R.Ā C., etĀ al. 2000, ApJ, 533, 682
- Carniani etĀ al. (2020) Carniani, S., Ferrara, A., Maiolino, R., etĀ al. 2020, MNRAS, 499, 5136
- Casey (2012) Casey, C.Ā M. 2012, MNRAS, 425, 3094
- Casey etĀ al. (2025) Casey, C.Ā M., Akins, H.Ā B., Finkelstein, S.Ā L., etĀ al. 2025, submitted to AAS journals, arXiv:2505.18873
- Casey etĀ al. (2024) Casey, C.Ā M., Akins, H.Ā B., Shuntov, M., etĀ al. 2024, ApJ, 965, 98
- Chabrier (2003) Chabrier, G. 2003, PASP, 115, 763
- Cowie etĀ al. (2017) Cowie, L.Ā L., Barger, A.Ā J., Hsu, L.Ā Y., etĀ al. 2017, ApJ, 837, 139
- da Cunha etĀ al. (2013) da Cunha, E., Groves, B., Walter, F., etĀ al. 2013, ApJ, 766, 13
- de Graaff etĀ al. (2025) de Graaff, A., Rix, H.-W., Naidu, R.Ā P., etĀ al. 2025, submitted to A&A, arXiv:2503.16600
- De Looze etĀ al. (2014) De Looze, I., Cormier, D., Lebouteiller, V., etĀ al. 2014, A&A, 568, A62
- De Rossi etĀ al. (2018) De Rossi, M.Ā E., Rieke, G.Ā H., Shivaei, I., Bromm, V., & Lyu, J. 2018, ApJ, 869, 4
- Decarli etĀ al. (2023) Decarli, R., Pensabene, A., Diaz-Santos, T., etĀ al. 2023, A&A, 673, A157
- DāEugenio etĀ al. (2025) DāEugenio, F., Cameron, A.Ā J., Scholtz, J., etĀ al. 2025, ApJS, 277, 4
- Egami etĀ al. (2023) Egami, E., Sun, F., Alberts, S., etĀ al. 2023, Complete NIRCam Grism Redshift Survey (CONGRESS), JWST Proposal. Cycle 2, ID. #3577
- Eisenstein etĀ al. (2023) Eisenstein, D.Ā J., Willott, C., Alberts, S., etĀ al. 2023, submitted to ApJS, arXiv:2306.02465
- Elbaz etĀ al. (2011) Elbaz, D., Dickinson, M., Hwang, H.Ā S., etĀ al. 2011, A&A, 533, A119
- Endsley etĀ al. (2022) Endsley, R., Stark, D.Ā P., Bouwens, R.Ā J., etĀ al. 2022, MNRAS, 517, 5642
- Enia etĀ al. (2022) Enia, A., Talia, M., Pozzi, F., etĀ al. 2022, ApJ, 927, 204
- Franco et al. (2018) Franco, M., Elbaz, D., Béthermin, M., et al. 2018, A&A, 620, A152
- Fritz etĀ al. (2006) Fritz, J., Franceschini, A., & Hatziminaoglou, E. 2006, MNRAS, 366, 767
- Fudamoto etĀ al. (2024) Fudamoto, Y., Inoue, A.Ā K., Coe, D., etĀ al. 2024, ApJ, 961, 71
- Fudamoto etĀ al. (2023) Fudamoto, Y., Inoue, A.Ā K., & Sugahara, Y. 2023, MNRAS, 521, 2962
- Fudamoto etĀ al. (2021) Fudamoto, Y., Oesch, P.Ā A., Schouws, S., etĀ al. 2021, Nature, 597, 489
- Fujimoto etĀ al. (2022) Fujimoto, S., Brammer, G.Ā B., Watson, D., etĀ al. 2022, Nature, 604, 261
- Gardner etĀ al. (2023) Gardner, J.Ā P., Mather, J.Ā C., Abbott, R., etĀ al. 2023, PASP, 135, 068001
- Gómez-Guijarro et al. (2022) Gómez-Guijarro, C., Elbaz, D., Xiao, M., et al. 2022, A&A, 658, A43
- Gómez-Guijarro et al. (2023) Gómez-Guijarro, C., Magnelli, B., Elbaz, D., et al. 2023, A&A, 677, A34
- Gordon etĀ al. (2010) Gordon, K.Ā D., Galliano, F., Hony, S., etĀ al. 2010, A&A, 518, L89
- Gottumukkala etĀ al. (2024) Gottumukkala, R., Barrufet, L., Oesch, P.Ā A., etĀ al. 2024, MNRAS, 530, 966
- Greene etĀ al. (2024) Greene, J.Ā E., Labbe, I., Goulding, A.Ā D., etĀ al. 2024, ApJ, 964, 39
- Gruppioni et al. (2020) Gruppioni, C., Béthermin, M., Loiacono, F., et al. 2020, A&A, 643, A8
- Harikane etĀ al. (2020) Harikane, Y., Ouchi, M., Inoue, A.Ā K., etĀ al. 2020, ApJ, 896, 93
- Harshan etĀ al. (2024) Harshan, A., Tripodi, R., Martis, N.Ā S., etĀ al. 2024, ApJ, 977, L36
- Herrera-Camus etĀ al. (2018) Herrera-Camus, R., Sturm, E., GraciĆ”-Carpio, J., etĀ al. 2018, ApJ, 861, 95
- Hildebrand (1983) Hildebrand, R.Ā H. 1983, QJRAS, 24, 267
- Hopkins etĀ al. (2010) Hopkins, P.Ā F., Murray, N., Quataert, E., & Thompson, T.Ā A. 2010, MNRAS, 401, L19
- Ji etĀ al. (2025) Ji, X., Maiolino, R., Ćbler, H., etĀ al. 2025, submitted to MNRAS, arXiv:2501.13082
- Kocevski etĀ al. (2025) Kocevski, D.Ā D., Finkelstein, S.Ā L., Barro, G., etĀ al. 2025, ApJ, 986, 126
- Kocevski etĀ al. (2023) Kocevski, D.Ā D., Onoue, M., Inayoshi, K., etĀ al. 2023, ApJ, 954, L4
- Kokorev etĀ al. (2024) Kokorev, V., Caputi, K.Ā I., Greene, J.Ā E., etĀ al. 2024, ApJ, 968, 38
- KovÔcs et al. (2006) KovÔcs, A., Chapman, S. C., Dowell, C. D., et al. 2006, ApJ, 650, 592
- Labbe etĀ al. (2025) Labbe, I., Greene, J.Ā E., Bezanson, R., etĀ al. 2025, ApJ, 978, 92
- Labbé et al. (2023) Labbé, I., van Dokkum, P., Nelson, E., et al. 2023, Nature, 616, 266
- Lagache etĀ al. (2018) Lagache, G., Cousin, M., & Chatzikos, M. 2018, A&A, 609, A130
- Laporte etĀ al. (2017) Laporte, N., Ellis, R.Ā S., Boone, F., etĀ al. 2017, ApJ, 837, L21
- Maiolino etĀ al. (2025) Maiolino, R., Risaliti, G., Signorini, M., etĀ al. 2025, MNRAS, 538, 1921
- Matthee etĀ al. (2024) Matthee, J., Naidu, R.Ā P., Brammer, G., etĀ al. 2024, ApJ, 963, 129
- McKinney etĀ al. (2023) McKinney, J., Pope, A., Kirkpatrick, A., etĀ al. 2023, ApJ, 955, 136
- McMullin etĀ al. (2007) McMullin, J.Ā P., Waters, B., Schiebel, D., Young, W., & Golap, K. 2007, in Astronomical Society of the Pacific Conference Series, Vol. 376, Astronomical Data Analysis Software and Systems XVI, ed. R.Ā A. Shaw, F.Ā Hill, & D.Ā J. Bell, 127
- Meyer etĀ al. (2024) Meyer, R.Ā A., Oesch, P.Ā A., Giovinazzo, E., etĀ al. 2024, MNRAS, 535, 1067
- Meyer etĀ al. (2025) Meyer, R.Ā A., Walter, F., Di Mascia, F., etĀ al. 2025, A&A, 695, L18
- Morishita etĀ al. (2024) Morishita, T., Stiavelli, M., Chary, R.-R., etĀ al. 2024, ApJ, 963, 9
- Naidu etĀ al. (2025) Naidu, R.Ā P., Matthee, J., Katz, H., etĀ al. 2025, arXiv e-prints, arXiv:2503.16596
- Oesch etĀ al. (2023) Oesch, P.Ā A., Brammer, G., Naidu, R.Ā P., etĀ al. 2023, MNRAS, 525, 2864
- Oke & Gunn (1983) Oke, J.Ā B. & Gunn, J.Ā E. 1983, ApJ, 266, 713
- Owen (2018) Owen, F.Ā N. 2018, ApJS, 235, 34
- Pallottini etĀ al. (2022) Pallottini, A., Ferrara, A., Gallerani, S., etĀ al. 2022, MNRAS, 513, 5621
- Pérez-GonzÔlez et al. (2023) Pérez-GonzÔlez, P. G., Barro, G., Annunziatella, M., et al. 2023, ApJ, 946, L16
- Pérez-GonzÔlez et al. (2024) Pérez-GonzÔlez, P. G., Barro, G., Rieke, G. H., et al. 2024, ApJ, 968, 4
- Planck Collaboration etĀ al. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., etĀ al. 2020, A&A, 641, A6
- Sacchi & Bogdan (2025) Sacchi, A. & Bogdan, A. 2025, submitted to ApJL, arXiv:2505.09669
- Salpeter (1955) Salpeter, E.Ā E. 1955, ApJ, 121, 161
- Schaerer et al. (2020) Schaerer, D., Ginolfi, M., Béthermin, M., et al. 2020, A&A, 643, A3
- Schaerer etĀ al. (2024) Schaerer, D., Marques-Chaves, R., Xiao, M., & Korber, D. 2024, A&A, 687, L11
- Schouws etĀ al. (2023) Schouws, S., Bouwens, R., Smit, R., etĀ al. 2023, ApJ, 954, 103
- Schreiber etĀ al. (2015) Schreiber, C., Pannella, M., Elbaz, D., etĀ al. 2015, A&A, 575, A74
- Setton etĀ al. (2025) Setton, D.Ā J., Greene, J.Ā E., Spilker, J.Ā S., etĀ al. 2025, submitted to ApJL, arXiv:2503.02059
- Shu etĀ al. (2022) Shu, X., Yang, L., Liu, D., etĀ al. 2022, ApJ, 926, 155
- Sommovigo etĀ al. (2022) Sommovigo, L., Ferrara, A., Pallottini, A., etĀ al. 2022, MNRAS, 513, 3122
- Stefanon et al. (2021) Stefanon, M., Labbé, I., Oesch, P. A., et al. 2021, ApJS, 257, 68
- Talia etĀ al. (2021) Talia, M., Cimatti, A., Giulietti, M., etĀ al. 2021, ApJ, 909, 23
- Tripodi etĀ al. (2023) Tripodi, R., Feruglio, C., Kemper, F., etĀ al. 2023, ApJ, 946, L45
- Tsukui etĀ al. (2023) Tsukui, T., Wisnioski, E., Krumholz, M.Ā R., & Battisti, A. 2023, MNRAS, 523, 4654
- Valentino etĀ al. (2023) Valentino, F., Brammer, G., Gould, K. M.Ā L., etĀ al. 2023, ApJ, 947, 20
- van der Vlugt etĀ al. (2023) van der Vlugt, D., Hodge, J.Ā A., Jin, S., etĀ al. 2023, ApJ, 951, 131
- Walter etĀ al. (2022) Walter, F., Neeleman, M., Decarli, R., etĀ al. 2022, ApJ, 927, 21
- Wang etĀ al. (2025) Wang, B., de Graaff, A., Davies, R.Ā L., etĀ al. 2025, ApJ, 984, 121
- Wang etĀ al. (2024) Wang, B., Leja, J., de Graaff, A., etĀ al. 2024, ApJ, 969, L13
- Wang etĀ al. (2016) Wang, T., Elbaz, D., Schreiber, C., etĀ al. 2016, ApJ, 816, 84
- Wang etĀ al. (2019) Wang, T., Schreiber, C., Elbaz, D., etĀ al. 2019, Nature, 572, 211
- Weibel etĀ al. (2024) Weibel, A., Oesch, P.Ā A., Barrufet, L., etĀ al. 2024, MNRAS, 533, 1808
- Whitaker etĀ al. (2019) Whitaker, K.Ā E., Ashas, M., Illingworth, G., etĀ al. 2019, ApJS, 244, 16
- Williams etĀ al. (2024) Williams, C.Ā C., Alberts, S., Ji, Z., etĀ al. 2024, ApJ, 968, 34
- Williams etĀ al. (2019) Williams, C.Ā C., Labbe, I., Spilker, J., etĀ al. 2019, ApJ, 884, 154
- Xiao etĀ al. (2024) Xiao, M., Oesch, P.Ā A., Elbaz, D., etĀ al. 2024, Nature, 635, 311
- Xiao et al. (2023) Xiao, M. Y., Elbaz, D., Gómez-Guijarro, C., et al. 2023, A&A, 672, A18
- Xue etĀ al. (2016) Xue, Y.Ā Q., Luo, B., Brandt, W.Ā N., etĀ al. 2016, ApJS, 224, 15
- Yue etĀ al. (2024) Yue, M., Eilers, A.-C., Ananna, T.Ā T., etĀ al. 2024, ApJ, 974, L26
- Zanella etĀ al. (2018) Zanella, A., Daddi, E., Magdis, G., etĀ al. 2018, Monthly Notices of the Royal Astronomical Society, 481, 1976
Appendix A Best-fit SED of ID2756 with K.